### v1

# Temperature Dependent Mechanical and Structural Properties of Uniaxially Strained Planar Graphene

Abstract
Using molecular dynamics simulations in a planar graphene sheet, we investigate the temperature dependence of its mechanical behavior under uniaxial tensile stress applied either along the armchair or the zigzag direction. Stress-strain curves are calculated for different temperatures and the corresponding dependence of various elastic parameters, like the Young modulus, the third-order elastic modulus, the tensile strength and failure strain, is presented. Fracture stress and strain, as well as the Young modulus, decrease almost linearly with temperature. The distributions of bond lengths and bond angles at different strains and temperatures are also discussed and approximate analytical expressions are presented. The latter describe accurately the numerically obtained distributions.
Keywords graphene  $\cdot$ molecular dynamics  $\cdot$ stress-strain response  $\cdot$ elastic properties  $\cdot$ bond length and angle distributions
1 Introduction
Since the discovery of graphene, there have been a number of investigations into its mechanical behavior. Despite the difficulty of applying controlled mechanical loads at the nanoscale, experimental studies have verified an exceptional value of stiffness and extremely high tensile strength ref-Lee-Science-2008 ; ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 , in accordance with corresponding theoretical predictions. There exist a number of related numerical investigations using molecular dynamics (MD) simulations with a variety of potential functions ref-Zakharchenko-PRL-2009 ; ref-Zhao-NanoLet-2009 ; ref-Neek-Amal-PRB-2010 ; ref-Shen-MatDes-2010 ; ref-Tsai-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-Wei-JCP-2011 ; ref-Dmitriev-CompMatSci-2012 ; ref-Kalosakas-JAP-2013 ; ref-Sgouros-2DMat-2016 ; ref-Souros-SciRep-2018 ; ref-Lebedeva-PhysE-2019 ; ref-Savin-MechofMat-2019 ; ref-Aparicio-Nanotech-2021 ; ref-deSousa-Intech-2021 ; ref-Kalosakas-Mat-2021 ; ref-Flores-PhysChem-2022 ; ref-Fthenakis-FrontChem-2022 ; ref-Pacheco-Sanjuan-Carbon-2023 ; ref-Sgouros-PRE-2024 , density functional theory ref-Kudin-PRB-2001 ; ref-Liu-PRB-2007 ; ref-Zeinalipour-Yazdi-JAP-2009 ; ref-Marianetti-PRL-2010 ; ref-Wagner-PRB-2011 ; ref-Mirnezhad-JTS-2012 ; ref-Shao-JCP-2012 , or other theoretical approaches including molecular mechanics ref-Lu-IJAM-2009 ; ref-Fthenakis-JoP-2017 ; ref-Genoese-FrontMat-2019 , and a combination of continuum elasticity theory with other methods ref-Cadelano-PRL-2009 ; ref-Samadikhah-CompMatSci-2012 ; ref-Wei-NanoLet-2013 .
The temperature dependence of various elastic properties of graphene has been examined using MD ref-Shen-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-deSousa-Intech-2021 , Monte Carlo atomistic simulations ref-Zakharchenko-PRL-2009 , density functional theory ref-Shao-JCP-2012 , or the asymptotic homogenization method ref-Tahani-AMM-2024 . Moreover, a few MD studies have investigated the variation of bond lengths and bond angles with uniaxial tensile loading ref-Zhao-NanoLet-2009 ; ref-Kalosakas-Mat-2022 . The latter one shows results from first principles methods too, and presents analytical expressions for the dependence of these structural parameters on strain ref-Kalosakas-Mat-2022 . The variation of bond lengths and angles with biaxial strain has been given in ref-Dheeraj-Nanotech-2020 . To our knowledge, the effects of temperature and strain on the distributions of bond lengths and bond angles have not yet been presented.
In the present work, we use MD simulations to study the behavior of planar graphene under uniaxial tensile load, considering the influence of temperature. In particular, we implement symplectic integration methods for simulating the system’s time evolution, which allow highly accurate computations for arbitrarily long times. We calculate stress-strain curves at various temperatures and from these results we estimate the variation of several elastic parameters with temperature. Furthermore, we compute bond lengths and bond angles of bulk graphene over a large time-window after thermal equilibrium has been reached, and subsequently analyze these results in order to get the dependence of the corresponding distributions on both stress and temperature. Finally, we present analytical expressions which closely match the numerically obtained distributions of bond lengths and angles. Thus, we describe the detailed dependence of bond lengths and bond angles in graphene on both the applied tensile stress and temperature.
The paper is organized as follows. In Section 2 we present the used force field, along with the numerical methods we implement. The results of our investigation are discussed in Section 3. In particular, the implementation of finite temperatures in our MD microcanonical simulations is discussed in Section 3.1, and thermal effects on graphene’s mechanical response under uniaxial tension are studied in Section 3.2. Then we present the distribution of bond lengths and bond angles in the planar sheet of $sp^{2}$ carbon atoms at various stresses and temperatures in Section 3.3, while analytical expressions for said distributions are determined in Section 3.3.1. Finally we conclude our findings in Section 4.
2 Model and numerical methods
We consider a two-dimensional (2D) model of graphene as a hexagonal lattice of carbon atoms within a plane. Figure 1 illustrates a part of this structure at equilibrium, where the distance between any two neighboring atoms is $r_{0}=1.42$ and the angle formed by three consecutive atoms is $\phi_{0}=2\pi/3$ rad. Furthermore, all carbon atoms have mass $m=12$ amu. In the orientation depicted in Figure 1, the top and bottom edges represent the "armchair edges", while the left and right edges correspond to the "zigzag edges". Furthermore, it is common to call "armchair direction" the horizontal direction in Figure 1, and "zigzag direction" the vertical one.
Figure 1: A schematic of the hexagonal graphene lattice containing $N=42$ atoms, arranged in $N_{I}=6$ columns and $N_{J}=7$ rows. Atoms in column $i$ and row $j$ are indicated in blue and orange, respectively, and the $(i,j)$th atom is colored in gray. The $A$ and $Z$ type bonds, and similarly the $\alpha$ and $\zeta$ type angles, are respectively indicated in red and green (see Section 3.3 for more details on these distinctions).
We refer to columns and rows within the graphene sheet, as indicated by the blue and orange colored atoms, respectively, in Figure 1. A lattice of size $N=N_{I}\times N_{J}$ consists of $N_{I}$ columns, which are indexed by $i$, and $N_{J}$ rows, which are indexed by $j$. The $(i,j)$th atom is indicated in gray in Figure 1. Thus, the configuration shown in this figure corresponds to a $N_{I}\times N_{J}=6\times 7$ lattice, totaling $N=42$ atoms. In our simulations we consider a lattice of $N_{I}=86$ columns and $N_{J}=87$ rows of atoms, resulting in a $N=86\times 87=7482$ total number of carbon atoms. This lattice size is sufficiently large to negate small-size effects ref-Kalosakas-JAP-2013 , allowing thus to represent results on bulk graphene, while being small enough to permit extensive numerical simulations within reasonable CPU times.
A Hamiltonian formalism is used to investigate the in-plane dynamics of the lattice in a similar fashion as that of Ref. ref-Hillebrand-Chaos-2020 . The atomistic force fields describing bond stretching and angle bending deformations are determined through fittings with relevant density functional energy computations ref-Kalosakas-JAP-2013 . In particular, the potential energy of a covalent bond between neighboring atoms at distance $r$ is given by the Morse expression,
$V_{M}(r)=D\left(e^{-a\left(r-r_{0}\right)}-1\right)^{2},$ (1)
where $D=5.7$ eV is the depth of the potential and $a=1.96$-1 is the inverse characteristic length scale of the potential. The angle bending energy term describing a bond angle $\phi$ formed by three consecutive atoms is
$V_{B}(\phi)=\frac{d}{2}\left(\phi-\phi_{0}\right)^{2}-\frac{d^{\prime}}{3}\left(\phi-\phi_{0}\right)^{3},$ (2)
where $d=7.0$ eV/rad2 and $d^{\prime}=4$ eV/rad3 are the quadratic and nonlinear coefficients of the potential, respectively.
The total energy of the system (i.e., the values of the model’s Hamiltonian $H$) is the sum of the above potential energy terms for all bond lengths of neighboring atoms and all bond angles between adjacent bonds, and the kinetic energy of each atom. Denoting the total potential energy at time $t$ by $E_{V}(t)$ and the total kinetic energy at $t$ by $E_{K}(t)$, the Hamiltonian
$H(t)=E_{K}(t)+E_{V}(t),$ (3)
is expressed through the $(x(t),y(t))$ positions and the corresponding conjugate momenta of all carbon atoms of the sheet. The time evolution of each atom’s position is governed by the system’s Hamilton’s equations of motion, which conserve the total energy (3).
To study the effects of uniaxial tensile load, a constant force $f$ is applied to all atoms on the appropriate edges ref-Kalosakas-JAP-2013 ; ref-Kalosakas-Mat-2022 : For stress/strain along the armchair direction, $\sigma_{a}$/$\epsilon^{a}$, the force $f$ is applied on the atoms of the zigzag edges, where on the opposite edges opposite forces, directed outwards, are applied. Similarly, for stress/strain along the zigzag direction, $\sigma_{z}$/$\epsilon^{z}$, the force $f$ is applied to the armchair edges, again with opposite forces on opposite edges. Tensile loading leads to additional terms in the Hamiltonian, concerning the relevant edge atoms where the force $f$ is applied. For constant forces, as in our case here, the conservation of the system’s total energy still holds.
In two-dimensional materials like graphene, the stress is given by the force per unit length. Taking into account the distance between successive atoms at the relevant edges where the force is applied, i.e. the atoms of an edge column (row) in Figure 1 concerning stress along the armchair (zigzag) direction, the following relations connect nominal stresses and forces
$\sigma_{a}=\frac{f}{r_{0}\sin(\phi_{0}/2)}\,\,\,\,\mbox{and}\,\,\,\,\sigma_{z}=\frac{f}{0.5\,r_{0}\left[1+\cos(\phi_{0}/2)\right]}.$ (4)
where $r_{0}$ and $\phi_{0}$ are the equilibrium values mentioned above.
To determine at zero temperature the relaxed state of the strained graphene for various applied stresses in any direction, a friction term proportional to the velocity of each atom is incorporated in the MD simulations, setting the friction proportionality coefficient to $\gamma=0.1$ ps-1 (see Ref. ref-Kalosakas-JAP-2013 ). Then, the fourth order Runge-Kutta numerical integration technique is used with an integration time step of $dt=0.005~{}t_{u}$, where $t_{u}=0.0102$ ps represents the time unit for our model. This time step ensures that the relative energy error $Err(t)=\left|H(t)-H(0)\right|/\left|H(0)\right|$ is below $10^{-7}$ in corresponding energy conserving simulations (without the friction term). However, we now simulate the dynamics of the dissipative version of the system until times $t_{f}=3\times 10^{3}~{}t_{u}$, when the total kinetic energy is practically zero ($E_{K}(t_{f})<10^{-16}$ eV). In this way we determine the relaxed equilibrium positions of the atoms for each considered stress $\sigma$.
Based on this equilibrium data of graphene subjected to tensile loads, we embark on the main phase of the numerical investigation: modeling the dynamics of the lattice for different values of stress $\sigma$, at various temperatures $T$, for a long enough time to allow deductions about the thermal equilibrium properties of the stressed material. For these numerical simulations we implement the symplectic integrator ABA864 ref-Blanes-ANM-2013 , with an integration time step $dt=0.06~{}t_{u}$, which results in $Err(t)<10^{-7}$ for all times. This particular integration scheme has been shown to perform very well in balancing computational speed and numerical accuracy for multidimensional Hamiltonian lattices ref-Danieli-MINE-2019 , and was successfully used for examining chaos in graphene ref-Hillebrand-Chaos-2020 .
The relaxed equilibrium positions that have previously been determined for the given value of stress $\sigma$ correspond to a graphene sheet being at zero temperature, without thermal fluctuations. In order to simulate the system at finite temperatures, following a suitable energy-temperature calibration (see subsection 3.1 below), we insert an additional energy density (average energy per site) $e_{N}$ on the relaxed $T=0$ K state. This additional energy is initially provided as solely potential energy, in the form of small random displacements of each atom from the relaxed zero temperature positions. Then these displacements are properly scaled in order to adjust the added energy density $e_{N}$ to the desired value. During evolution the initial potential energy gets shared into kinetic and potential energy and finally the system equilibrates.
In general, for the numerical results presented below we consider 10-20 different individual realizations of the randomly added initial energy, but we have selectively checked the robustness of the data when more realizations are used. For each realization we calculate the temporal evolution of the various quantities of interest, and then compute the average of these time-series over the different realizations in order to obtain the time dependence of the considered quantities for the ensemble. We denote the averaged quantity over the different realizations with angled brackets, e.g. $\left\langle M(t)\right\rangle$ for the measurement of the quantity $M(t)$. We may further determine the average of a thermally equilibrated quantity over time. In such a case we average both over initial realizations and over time intervals, and we denote the computed average by using both an overline and angled brackets, e.g. $\overline{\left\langle M\right\rangle}$ for a variable $M$ at thermal equilibrium.
At finite temperatures the size of graphene sheets exhibits oscillations around their $T=0$ K relaxed configurations due to the thermal energy of the system (discussed further in Section 3.2 below). In order to collect data over sufficiently many such sheet oscillations, we follow the system’s time evolution up to $4\times 10^{3}~{}t_{u}$. Thus, the recording window for all subsequent measurements is from $1\times 10^{3}$ to $4\times 10^{3}~{}t_{u}$, totaling $3000~{}t_{u}$. We have checked the insensitivity of the obtained results on the length of the recording window by testing longer time windows.
3 Results and Discussion
3.1 Temperature calibration
After inserting an energy density $e_{N}$ in the strained lattice, we observe that initially $T$ increases with time from its zero starting value and then following some relatively large fluctuations the system is settled at thermal equilibrium after at most $10^{3}~{}t_{u}$. The temperature $T(t)$ as a function of time is computed in our microcanonical MD simulations through the energy equipartition relation
$T(t)=\frac{E_{K}(t)}{N\,k_{B}}$ (5)
where $k_{B}$ is the Boltzmann constant.
In order to test that thermal equilibrium has been reached, we compare the mean and standard deviation of the fluctuating temperature over various time windows. Before achieving thermal equilibrium the standard deviation of the time-averaged $T$ is relatively large and also changes depending on the time window. When thermal equilibrium is reached the temperature fluctuations and the standard deviation are consistently small. The mean temperature at thermal equilibrium is calculated by averaging over both the individual realizations and the recording time-window. We denote this average value $\overline{\left\langle T\right\rangle}$ by $T_{ave}$. In this case, the standard deviation of the measured values is computed using all data points over realizations and time.
The relationship between the additional energy density $e_{N}$ above the relaxed equilibrium loaded structures and the averaged temperature $T_{ave}$ is linear for all cases examined here, corresponding to temperatures up to 700 K. One representative case is shown in Figure 2. The resulting slopes of the data fittings are very close for all values of stress $\sigma$ (a difference in the computed values is observed only in the fourth significant digit) and they are slightly above $2k_{B}$ due to the nonlinearities of the potential energy. For finite loads, the slope slightly increases with the amount of stress. For a given value of stress $\sigma$, we use the slope of the $e_{N}$ versus $T_{ave}$ fitting in order to control temperature (within a 1% accuracy) in our investigation. In particular, we are setting the amount of the added energy density $e_{N}$ according to the desired temperature value.
Figure 2: Relation between the average temperature $T_{ave}$ at thermal equilibrium, evaluated through averaging over both time and the different realizations, and the energy density $e_{N}$ above the relaxed $T=0$ K graphene structure subjected to uniaxial tensile stress $\sigma_{z}=2.16$ eV/Å2 along the zigzag direction. One standard deviation of the $T_{ave}$ measurements is indicated by blue error-bars. The slope of the linear fitting of the presented data points (indicated by the gray line) equals to $1.74\times 10^{-4}$ eV/K.
In order to avoid potential edge effects and represent the behavior of bulk graphene, we collect data from the central region of the lattice to investigate the system’s elastic and structural properties discussed below. In particular, this sub-lattice has an analogous geometry to the larger structure, with number of columns $n_{I}=44$ and number of rows $n_{J}=45$.
3.2 Mechanical response
In the stress-controlled numerical implementation used here to examine the mechanical response of graphene, we need to compute the resulting strain due to the fixed force applied at the appropriate edge atoms. The uniaxial strain $\epsilon$ is obtained through the strain of the central row of the graphene sheet when the stress is applied along the armchair direction, while it is calculated by the average strain on the two central columns of the sheet in case of stress along the zigzag direction (see Figure 1). By $\epsilon_{T}$ is indicated the strain corresponding to temperature $T$.
In the zero temperature case, $T=0$ K, the uniaxial strain $\epsilon_{0}$ in the bulk is determined through the relative change of the length of the central row (two central columns) of graphene subjected to the applied stress along the armchair (zigzag) direction, with respect to the length of the central row (columns) at the unstrained equilibrium configuration shown in Figure 1, where we take these measurements in the central region of the lattice. We clarify that for any length computations discussed here, the horizontal (vertical) length is measured as the difference of the $x$ ($y$) coordinates of the considered edge atoms. The stress-strain response is obtained in this way at 0 K
When the temperature of the system is raised at finite values by adding energy to the equilibrated graphene, the lattice stretches and compresses in an oscillatory manner. The details of these oscillations depend on the temperature and the applied stress and will be discussed in a future publication. In this case one has to take into account that the strain measurement $\epsilon_{T}(t)$ is now exhibiting temporal oscillations. Since we consider 10-20 realizations of the random initial energy distribution, we register the average, over these realizations, strain in time $\left\langle\epsilon_{T}(t)\right\rangle$, noting that the aforementioned oscillations are in-phase in the different realizations. For evaluating the strain $\epsilon_{T}(t)$, the reference length $\ell_{T}^{r}$ corresponding to zero stress $\sigma=0$ at temperature $T$ is obtained by calculating the average, over realizations and time, of the length of the central row or columns of the sheet in the absence of any load. Then, when a stress is applied the time evolution of strain in a particular realization is computed as
$\epsilon_{T}(t)=\frac{\ell(t)-\ell_{T}^{r}}{\ell_{T}^{r}}.$ (6)
where $\ell(t)$ is the length in time of the central row or the average length of the two central columns depending on the direction of the applied uniaxial load.
In Figure 3 we highlight the behavior of $\left\langle{\epsilon_{T}}(t)\right\rangle$ for various values of stress $\sigma$ along the zigzag direction, at three distinct temperatures $T$ shown by different colors. An increase in temperature leads to an increase in the amplitude of strain oscillations as well as in an increase in the average strain. The latter one is obtained as the average over both realizations and time, $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ and it is indicated by the dashed horizontal lines of different colors depending on the temperature in Figure 3. The average strain $\overline{\left\langle{\epsilon_{700}}\right\rangle}$ (red horizontal dashed lines) about which the $T=700$ K curves oscillate are higher than the $\overline{\left\langle{\epsilon_{100}}\right\rangle}$ (blue horizontal dashed lines) of the $T=100$ K curves in all cases of different stress. However, these differences are larger on absolute values for larger stresses.
Figure 3: Time evolution of the average (over individual realizations) strain $\left\langle\epsilon_{T}(t)\right\rangle$, Equation (6), when a stress (a) $\sigma=0.188$ eV/Å2, (b) $\sigma=1.03$ eV/Å2, and (c) $\sigma=1.97$ eV/Å2, along the zigzag direction is applied, for different temperatures $T=100$ K (blue curves), $T=400$ K (green curves), and $T=700$ K (red curves). The average (over realizations and time) strain $\overline{\left\langle\epsilon_{T}\right\rangle}$, for each temperature, is indicated by the horizontal dashed lines of the same color in each panel.
Calculating the average strain $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ as mentioned above, the mechanical response of planar graphene at different temperatures is obtained. Stress-strain curves for uniaxial tensile loads along the armchair and zigzag directions are presented for various temperatures in Figure 4. Despite the small differences, one can see for larger stresses that the average strain is a bit further to the right for the higher temperature cases. The error-bars indicate the standard deviation of the average strain measurement. As one can also deduce from Figure 3, the standard deviation is higher for higher temperatures. We further note that these strain measurements have been tested for larger numbers of realizations or longer time windows. In particular, the measured strain differs in the 3rd significant digit at most, when increasing the number of realizations or doubling the length of the time window.
Figure 4: Stress-strain response of planar graphene for uniaxial loads along the (a) armchair (b) zigzag direction, for different temperatures as indicated in the legend. Filled circles indicate the measured average strain for each given stress. Solid curves represent fittings of these data points with Equation (7). For $T\neq 0$ K the strain is given as the average over time and realizations, $\overline{\left\langle\epsilon_{T}\right\rangle}$, and the error-bars correspond to one standard deviation.
Since in Figure 4 the strain is measured with respect to the average length $\ell_{T}^{r}$ due to thermally induced oscillations, the stress-strain curves pass from the origin of Figure 4, as expected. From Figure 4, we see that the temperature has a relatively small effect in the stress-strain response, at least for the values of $T$ considered here, apart from the significant reduction of the fracture point. For small stresses the achieved strain is practically the same for the two directions of applied stress, while the strong directional dependence at large stresses has already been well established in previous investigations ref-Liu-PRB-2007 ; ref-Zhao-NanoLet-2009 ; ref-Kalosakas-JAP-2013 ; ref-deSousa-Intech-2021 .
The stress-strain response can be described by the nonlinear relation
$\sigma=E_{2D}\cdot\epsilon+D_{2D}\cdot\epsilon^{2},$ (7)
where $\sigma$ is the applied uniaxial stress, $\epsilon$ the corresponding strain, $E_{2D}$ is the 2D Young’s modulus and $D_{2D}$ is the 2D third-order elastic modulus. For each temperature examined and both directions of applied stress, we fit the data presented in Figure 4 with Equation (7) to determine graphene’s elastic moduli. The computed values of $E_{2D}$ and $D_{2D}$ are plotted in Figure 5 as a function of temperature, for applied stress in either the armchair (red points) or the zigzag (blue points) direction. The error-bars on these points indicate one standard deviation of the fitted parameters under the observed covariance of the fit. Furthermore, linear fittings of these data points are indicated by the dashed red (dotted blue) line for stress in the armchair (zigzag) direction.
Figure 5: Temperature dependence of (a) the Young’s modulus $E_{2D}$, and (b) the third-order elastic modulus $D_{2D}$, for applied stress along the armchair (red points) or the zigzag (blue points) direction, evaluated through fittings of the data of Figure 4 with Equation (7).
Figure 5 shows that both elastic moduli exhibit an almost linear dependence on temperature. The Young’s modulus $E_{2D}$ [Figure 5(a)] appears to decrease linearly with increasing temperature, albeit only by a relatively small amount, which is consistent with other results in the literature ref-Shen-MatDes-2010 ; ref-Shao-JCP-2012 ; ref-Tahani-AMM-2024 . In particular, the linear fitting of these data for stress along the armchair direction leads to a variation of $E_{2D}(T)$ with a slope $-8.1\times 10^{-3}$ (N/m)/K, while for stress along the zigzag direction the slope is $-4.37\times 10^{-3}$ (N/m)/K. The decrease of Young’s modulus with temperature is often given in the literature as a percentage change over the investigated range of temperatures. In our case the Young’s modulus decreases by $1.8\%$ ($0.94\%$) for stress along the armchair (zigzag) direction over the temperature range from 0 K to 700 K. In ref-Shen-MatDes-2010 , MD is used for investigations over the range from 300 K to 700 K, and $E_{2D}$ is found to decrease by 5.1%, 1.9%, or 1.3% (decrease by 3.5%, 2.6%, or 1.3%) for strain in the armchair (zigzag) direction, where the three different measurements are for three different aspect ratios of the graphene lattice. In ref-Shao-JCP-2012 density functional theory is used to determine that $E_{2D}$ decreases by 2.2% over the range from 0 K to 1000 K, with no directional dependence indicated. In ref-Tahani-AMM-2024 this reduction in $E_{2D}$ is found to be between 3% and 4% (no directional dependence indicated) for the temperature ranging from 0 K to 1600 K, where the varying reduction depends on the exact assumptions made for the asymptotic homogenization model, which affects the value at $T=0$ K.
We observe from Figure 5(b) that the $D_{2D}$ values are consistently higher for strain in the zigzag direction, than for the other direction. This is congruent with the fact that the graphene sheet is more resistant to stress along the zigzag direction. Since there are no bonds parallel to the direction of strain when forces are applied along the zigzag direction, the sheet can tolerate larger stresses. When stress is along the armchair direction, one third of all the bonds are parallel to the direction of strain, and hence these bonds exhibit maximal stretching in the sheet, leading to higher strain for the same stress, and hence the lower $D_{2D}$ values. This higher strain per stress due to the particular direction can be seen when comparing the panels of Figure 4, where the curves for stress in the armchair direction lie further to the right than when the stress is applied in the zigzag direction, i.e., indicating higher strains for similar stresses. Bonds that are not parallel to the direction of strain also stretch out, but due to their orientation, more stress is required to stretch them as significantly as the parallel bonds are stretched. Regarding the temperature variation of $D_{2D}$, different trends are exhibited when the stress is along the zigzag or the armchair direction. A linear fitting of the $D_{2D}(T)$ data points results in a slope $+4.95\times 10^{-2}$ (N/m)/K for strain in the armchair direction and $-1.01\times 10^{-2}$ (N/m)/K for strain in the zigzag direction. The value of $D_{2D}$ increases by $6.3\%$ (decreases by $1.4\%$) for strain in the armchair (zigzag) direction over the temperature range from 0 K to 700 K.
Furthermore, we estimate the graphene’s tensile strength $\sigma_{f}$ and failure strain $\epsilon_{f}$, for different temperatures $T$. The former one is obtained by the highest tested value of $\sigma$ which does not lead to failure of the graphene sheet. Its error-bar is provided by the step we use in the increment of the tested $\sigma$ values, which are evenly spaced. These results are presented in Figure 6(a), where an almost linear decrease of the fracture stress with temperature is shown. A linear fitting of these data points is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The slope of the linear fitting of the $\sigma_{f}(T)$ data is $-8.37\times 10^{-3}$ (N/m)/K for stress in the armchair direction and $-1.45\times 10^{-2}$ (N/m)/K for stress in the zigzag direction. Such a linear dependence of the tensile strength on temperature is in accordance with existing results ref-Zhao-JAP-2010 . In particular, we can estimate that the reported tensile strength per temperature has a slope of $-8.57\times 10^{-3}$ (N/m)/K in this aforementioned MD investigation of a 3936 atom graphene lattice.
Figure 6: Temperature dependence of (a) the tensile strength $\sigma_{f}$ and (b) the associated failure strain $\epsilon_{f}$ of graphene. Straight lines represent linear fittings.
We have also estimated the failure strain, $\epsilon_{f}$, at different temperatures, by solving for $\epsilon_{f}$ the Equation (7). To this end, the known value of $\sigma_{f}$ as well as the fitted values of $E_{2D}$ and $D_{2D}$ describing the stress-strain curve at the given temperature, have been used. In this case the error-bars are determined by converting the corresponding extreme values of stress, $\sigma_{f}\pm\Delta\sigma_{f}$, to strain [via Equation (7)], and then choosing the maximum absolute difference from $\epsilon_{f}$. These results are shown in Figure 6(b) and again a linear fitting of the data is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The linear fitting of the $\epsilon_{f}(T)$ data points leads to a slope $-4.71\times 10^{-3}$ strain/K for stress in the armchair direction and $-1.60\times 10^{-2}$ strain/K for stress in the zigzag direction. It is reasonable that higher temperatures cause failure at lower stresses/strains, since the maximum deformation of the graphene lattice under a fixed load increases with temperature, as shown in Figure 3. Therefore the sheet breaks more easily at higher temperatures. Moreover, the known property that graphene can tolerate higher loads along the zigzag direction is clear by the results in Figure 6, where the values of $\sigma_{f}$ and subsequently $\epsilon_{f}$ are consistently lower for stress in the armchair direction (red data points) as compared to stress applied in the zigzag direction (blue data points).
3.3 Bond length and bond angle distributions
In order to analyze the effects of temperature and stress on the distributions of the lengths and angles of the bonds, we first distinguish the two types of bond lengths, which we denote $A$ and $Z$, and the two types of angles, which we denote by $\alpha$ and $\zeta$, as illustrated in Figure 1. The $A$ bonds are along the armchair direction (equivalently, perpendicular to the zigzag direction). The $Z$ bonds alternate symmetrically along the zigzag direction and both exhibit identical deformations at 0 K when a uniaxial stress is applied along a high symmetry direction, such as the zigzag or the armchair direction. The angles $\alpha$ and $\zeta$ represent the bond angles formed between two $Z$ bonds and between an $A$ and a $Z$ bond, respectively. They respond always differently under an applied stress, due to the geometry of the system and the constraint $\alpha+2\zeta=2\pi$.
When a load is applied at zero temperature, there is no variability in any type of bond length and angle due to the static nature of the strained sheet at 0 K. Approximate expressions for the strain dependence of bond lengths $A$ and $Z$ and angles $\alpha$ and $\zeta$ were determined in Ref. ref-Kalosakas-Mat-2022 . Indicating by the indices $a$ or $z$ a load applied along the armchair or zigzag direction, respectively, these expressions read
$A_{a}$ $=1.42+0.011\,\epsilon_{0}+0.00024\,\epsilon^{2}_{0},$ (8)
$Z_{a}$ $=1.42+0.0031\,\epsilon_{0}-0.000046\,\epsilon_{0}^{2},$ (9)
and
$\alpha_{a}$ $=120^{\circ}-0.83\,\epsilon_{0}+0.020\,\epsilon^{2}_{0},$ (10)
$\zeta_{a}$ $=120^{\circ}+0.41\,\epsilon_{0}-0.010\,\epsilon^{2}_{0},$ (11)
while for stress along the zigzag direction
$A_{z}$ $=1.42,$ (12)
$Z_{z}$ $=1.42+0.0088\,\epsilon_{0}+0.000080\,\epsilon_{0}^{2},$ (13)
and
$\alpha_{z}$ $=120^{\circ}+0.80\,\epsilon_{0}-0.013\,\epsilon^{2}_{0}$ (14)
$\zeta_{z}$ $=120^{\circ}-0.40\,\epsilon_{0}+0.0064\,\epsilon^{2}_{0}.$ (15)
In Equations (8), (9), (12) and (13), the constants have units Å, the coefficients of $\epsilon_{0}$ have units Å/(%strain), and the coefficients of $\epsilon_{0}^{2}$ have units Å/(%strain)2. Similarly, in Equations (10), (11), (14) and (15) the constants have units of deg, the coefficients of $\epsilon_{0}$ have units deg/(%strain), and the coefficients of $\epsilon_{0}^{2}$ have units of deg/(%strain)2.
The distributions of all the bond lengths and angles in the bulk graphene sheet at $T=0$ K are given by double singular peaks at the locations provided by the above pairs of relations for the bond lengths and angles, respectively, for different values of the applied uniaxial strain.
In order to reveal the influence of temperature on the bond lengths and angles distributions, we register all the bond length and angle values during the system’s evolution in our measurement window and obtain normalized distributions for different amounts of stress/stain at various temperatures. In particular, we create a distribution for each realization by allocating all the measured bond lengths (angles) into fine-grained bins of width $3.5\times 10^{-3}$ ($0.004\,\frac{180^{\circ}}{\pi}$). We normalize the resulting distributions and then average them over the different realizations to obtain the final distribution for each case. It is worth noting that the size of the error-bars, indicating one standard deviation of this averaging computation over the different realizations, are negligible, and hence not included in the plots of the distributions presented below. We emphasize again that we consider the central region of the sheet for collecting our data, as mentioned at the end of Section 3.1.
In Figures 7(a) - (d) we show the normalized bond length distributions for increasing values of stress applied along the armchair direction. In Figures 7(e) - (h), the applied load is along the zigzag direction and the stress increases from (e) to (h) too. When there is no stress, $\sigma=0$, at finite temperatures the distributions (not shown in this figure) are simply normal distributions with the variance linearly increasing with temperature (see Section 3.3.1 below). Increased temperature leads to larger fluctuations in the lattice, resulting in a wider spread of the observed bond lengths. For the smaller values of stress presented in Figures 7(a) and (e), there is a slight skewing of the distributions, but as the stress is increased the single peak splits into two peaks which are gradually separated more and more as can be clearly seen at the lower temperatures. However, the increase of temperature leads to the merging of these two peaks due to their broadening. The centers of the peaks correspond to the zero-temperature values of the two types of bond lengths for each different direction of the applied stress as given in Equations (8) - (9) and (12) - (13).
Figure 7: Normalized bond length distributions in graphene, for increasing applied stress (left to right), along the armchair (top row) and the zigzag (bottom row) direction, at different temperatures $T$ as indicated in the legend. The stresses $\sigma$ in the armchair direction are (a) 0.569 eV/Å2, (b) 0.895 eV/Å2, (c) 1.22 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.563 eV/Å2, (f) 1.13 eV/Å2, (g) 1.69 eV/Å2, and (h) 2.16 eV/Å2.
Since there are twice as many $Z$ type bond lengths as $A$ types, the highest peak in each distribution in Figure 7 is mostly encompassing the lengths of the $Z$ type bonds. Thus, we can see that for stress applied along the armchair direction [Figures 7(a) - (d)], it is the $A$ bonds which achieve greater lengths (the lower peak, further to the right) while the $Z$ bonds exhibit a smaller extension. In contrast, for stress applied along the zigzag direction [Figures 7(e) - (h)], the $Z$ bonds achieve greater lengths (the taller peak is to the right in the distributions) while the centers of the smaller peaks remain near $r_{0}=1.42$, in accordance with Equation (12) and the corresponding broadening due to thermal effects. The fact that all bonds stretch for stress applied along the armchair direction, but only the $Z$ type bonds (two-thirds of all the considered bonds) stretch for a load along the other direction ref-Kalosakas-Mat-2022 , explains why the gap between the two peaks is more pronounced for stress applied in the zigzag direction.
In Figure 8 we present similar results as in Figure 7, but for the distribution of bond angles. Again we note the peak splitting due to increased stress, while increasing the temperature leads to the broadening and merging of these peaks. At zero strain normal distributions centered about the equilibrium angle of $\phi_{0}=120^{\circ}$ are obtained, with a variance increasing with temperature (see Section 3.3.1). The highest peak in the bond angle distributions corresponds to the $\zeta$ angle, since there are twice as many $\zeta$ angles as $\alpha$ angles. For stress along the armchair direction [Figures 8(a) - (d)], the $\alpha$ angles decrease, while the $\zeta$ angles increase. Reverse is the situation when the stress is applied along the zigzag direction [Figures 8(e) - (h)]. We again note that the peaks of the distributions are centered about the zero temperature $\alpha$ and $\zeta$ values, as given in Equations (10) - (11) and (14) - (15), depending on the direction of the applied stress.
Figure 8: Normalized bond angle distributions in planar graphene, for increasing stress (left to right), applied along the armchair (top row) and in the zigzag (bottom row) direction, at various temperatures $T$ as indicated in the legend. The values of stress $\sigma$ along the armchair direction are (a) 0.244 eV/Å2, (b) 0.651 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while in the zigzag direction are (e) 0.282 eV/Å2, (f) 0.939 eV/Å2, (g) 1.60 eV/Å2, and (h) 2.16 eV/Å2.
3.3.1 Analytical expressions of bond length and bond angle distributions
We now present approximate analytical expressions for the bond length and angle distributions, as those shown in Figures 7 and 8, in order to describe the dependence of graphene’s structural properties on stress and temperature. Based on the results discussed in the previous Subsection, we note that the obtained distributions appear to approximately be given through the combination of two normal distributions, where the means of these normal distributions correspond to the two types of bond lengths, or angles, found for each stress at zero temperature. The variance of these normal distributions is induced by thermal fluctuations, while the difference in peak heights is related to the fact that there exist double as many of one type of bond length (or angle) as the other.
As there exist approximate expressions available for the equilibrium bond lengths and bond angles as a function of the applied strain at $T=0$ K [see Equations (8) - (15)], it remains to determine the explicit dependence of variance on temperature. This will be obtained by numerically evaluating the effects of temperature on the normal distributions of the bond lengths and bond angles, at the unstrained graphene sheet. We can compare the results of these calculations with analytical estimates of the variance through the Boltzmann distribution, using a second-order approximation on the relevant potential energy terms describing bond stretching and angle bending.
Performing a Gaussian curve fitting procedure to the numerically obtained distributions of the bond lengths and bond angles at zero applied stress for various temperatures, shown in Figures 9(a) and (b), respectively, we compute the corresponding variances and mean values. The dependence of these variances on temperature are presented by filled circles in Figures 9(c) and (d) for the bond length and angle distributions, respectively. Solid lines in the latter plots denote a linear fitting of the data. It is worth noting that the mean of the bond length distribution slightly increases with temperature too. This is due to the thermal expansion of the 2D lattice. However, incorporating this small variation of the mean value with temperature does not practically affect the results discussed below. The mean of the bond angle distribution does not change with temperature, as expected.
Figure 9: Normalized distributions of (a) bond lengths and (b) bond angles in bulk graphene for $\sigma=0$, at different temperatures $T$ as indicated in the legend. (c) data points show the dependence on $T$ of the computed variance $\Sigma_{M}^{2}$ of the bond length distributions presented in (a). Similarly, data points in (d) denote the variance $\Sigma_{B}^{2}$ of the bond angle distributions shown in (b) versus temperature. Solid lines in (c) and (d) indicate linear fitting of the corresponding data, see Equations (16) and (17), while dotted lines denote the analytical approximating expression of Equations (19) and (20).
The numerically determined linear fitting of the variance $\Sigma^{2}_{M}$ of bond length distributions with the temperature $T$ is
$\Sigma^{2}_{M}(T)=C_{M}\,T,$ (16)
with $C_{M}=1.66\times 10^{-6}$2/K [solid line in Figure 9(c)]. Similarly, the numerically found variances $\Sigma^{2}_{B}$ for the bond angle distributions are well described by
$\Sigma^{2}_{B}(T)=C_{B}\,T,$ (17)
with $C_{B}=1.02\times 10^{-2}$ deg2/K [solid line in Figure 9(d)].
The linear dependence of these variances on temperature can be derived by the Boltzmann distribution when a quadratic approximation of the corresponding potential energy is considered. In particular, by the second derivative of the Morse potential of Equation (1),
$V_{M}^{\prime\prime}(r)=-2Da^{2}e^{-2a(r-r_{0})}\left(e^{a(r-r_{0})}-2\right),$ (18)
the second order approximation of the bond stretching energy term about $r=r_{0}$ reads $V_{M}^{lin}=a^{2}D(r-r_{0})^{2}$. Using this approximation, the corresponding Boltzmann distribution $\exp\left(\frac{-a^{2}D(r-r_{0})^{2}}{k_{B}T}\right)$ results in a normal bond length distribution of the form $\exp\left(\frac{-(r-r_{0})^{2}}{2\Sigma_{M}^{2}}\right)$ centered about the mean $r_{0}$ with a variance
$\Sigma^{2}_{M}(T)=\frac{k_{B}}{2a^{2}D}\,T,$ (19)
which gives $\Sigma^{2}_{M}(T)=1.97\times 10^{-6}\,T$, where the coefficient of $T$ has units of Å2/K.
Following a similar approach to estimate the variance of the bond angle distributions for different temperatures, we consider the second-order approximation of the potential $V_{B}$, Equation (2), about $\phi=\phi_{0}=120^{\circ}$, given by $V_{B}^{lin}=\frac{d}{2}\frac{\pi^{2}}{180^{2}}(\phi-\phi_{0})^{2}$ (when angles are measured in degrees). Note that due to the constrains in the sums of the bond angles around a particular atom and within hexagonal rings, just one angle can not be varied alone, but at least three other angles also change. Thus, we incorporate a factor 4 into the obtained energy approximation in the Boltzmann distribution, and eventually find that the the variance (in degrees), about the mean $\phi_{0}$, of the bond angle distribution is
$\Sigma^{2}_{B}(T)=\frac{k_{B}180^{2}}{4\pi^{2}d}\,T,$ (20)
which results in $\Sigma^{2}_{B}(T)=1.01\times 10^{-2}\,T$, where the coefficient of $T$ has units of deg2/K.
Dotted lines in Figures 9(c) and (d) correspond to the analytical expressions of Equations (19) and (20), respectively. We can see from Figure 9(c) that the analytically predicted slope of Equation (19) is somehow larger than the corresponding numerical value (the relative difference is less than 20%). Concerning the variance of the bond angle distributions, Figure 9(d) shows an excellent agreement between the analytically and numerically obtained slopes, exhibiting a relative difference of less than 1%. The reason for the quantitative disparity between the analytical prediction and numerical determination of the slope in the linear temperature dependence of the variance of the bond length distributions, but not for the angle distributions, may be due to the fact that the second order approximation of the angle bending potential $V_{B}(\phi)$ of Equation (2) is valid for a wide range of angles (see Figure 2 in Ref. ref-Kalosakas-JAP-2013 ). On the other hand, due to the highly anharmonic nature of the Morse potential $V_{M}(r)$, Equation (1), in the same energy scales (see Figure 1 in Ref. ref-Kalosakas-JAP-2013 ), the second order approximation about $r_{0}$ is only valid very close to $r_{0}$. Thus for small temperatures, for which the achieved bond lengths are all nearby $r_{0}$, the analytical and numerical predictions in Figure 9(c) are quite close. Nevertheless, as the temperature $T$ increases, much larger bond lengths come into play and hence the analytical prediction of the variances significantly deviates from the numerical results.
Combining now the numerically determined variances for different temperatures and the known bond length and angle mean values as a function of the applied stress/strain, analytical approximate expressions for the bond length and angle distributions can be derived. Regarding the bond length distributions, an additional issue should be taken into account when the numerically determined variances from Equation (16) will be used. In particular, the relation between the variance and the temperature should be scaled according to the behavior of $V^{\prime\prime}_{M}$, Equation (18), at the mean of the corresponding peak of the distribution, since bond lengths even further away from $r_{0}$ are encountered once stress is applied to the system. Given that analytically the variance equates to $k_{B}T/V_{M}^{\prime\prime}(r_{0})$ close to $r=r_{0}$, we multiply the numerically determined variance from Equation (16) by the scaling function
$F(r)=\frac{V_{M}^{\prime\prime}(r_{0})}{V_{M}^{\prime\prime}(r)},$ (21)
where $r$ is the known mean of the peak of interest in the distribution.
As a result, the bond length distribution for a given applied stress/strain and temperature $T$ can be approximated by the relation
$P_{M}=\frac{1}{\sqrt{2\pi\,C_{M}F\left(A\right)\,T}}\exp\left(-\frac{(r-A)^{2}}{2\,C_{M}F\left(A\right)\,T}\right)+\frac{2}{\sqrt{2\pi\,C_{M}F\left(Z\right)\,T}}\exp\left(-\frac{(r-Z)^{2}}{2\,C_{M}F\left(Z\right)\,T}\right),$ (22)
where $A$ and $Z$ are functions of the applied stress/strain, determined in Equations (8) - (9) and Equations (12) - (13), respectively, $C_{M}$ is given in Equation (16), and $F(r)$ is provided by Equation (21). Note that the quantities $A$ and $Z$ are provided by the corresponding zero temperature relations in Equations (8), (9), (12), and (13) as a function of strain $\epsilon$. If they are needed as a function of stress $\sigma$, the stress-strain relation of Equation (7) should be used to change the variable of the applied load.
For the bond angle distribution, the subtlety mentioned above concerning the scaling function is not needed since the second derivative of the angle bending potential, Equation (2), is everywhere the same regardless of the angle value at the peak of the distribution. Therefore, the angle bending distributions can be approximated by the expression
$P_{B}=\frac{1}{\sqrt{2\pi\,C_{B}\,T}}\left[\exp\left(\frac{-(\phi-\alpha)^{2}}{2\,C_{B}\,T}\right)+2\,\exp\left(\frac{-(\phi-\zeta)^{2}}{2\,C_{B}\,T}\right)\right],$ (23)
where $\alpha$ and $\zeta$ are determined by the applied stress/strain from Equations (10), (11), (14), (15), and Equation (7) if necessary, while $C_{B}$ is given in (17).
Points in Figure 10 present the numerically computed bond length distributions at various applied stresses and temperatures, while the solid lines correspond to the normalized curve $P_{M}$ from Equation (22). Figure 11 contains similar results, but for the bond angle distributions. From the plots of Figure 11 we see that the analytical expression $P_{B}$, Equation (23), describes the data very well, apart from small discrepancies at the heights of the taller peak at the lower $T$ depicted and for the smaller values of stress. On the other hand, the analytical expression $P_{M}$ from Equation (22) in Figure 10 deviates from the numerically obtained distributions for the longer-bond peak at the larger values of applied stresses [Figures 10(d), (g) and (h)]. Nonetheless, both expressions of Equations (22) and (23) provide a useful analytical description of the underlying structural properties of the strained graphene at finite temperatures.
Figure 10: Bond length distributions for selected temperatures $T$ (as shown in the legend), for stress along the armchair direction (left column) and the zigzag direction (right column), for increasing loads from top to bottom. In particular, the stresses $\sigma$ in the armchair direction are (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2. The stresses along the zigzag direction are (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions of Equation (22) are indicated by the solid curves.
Figure 11: Bond angle distributions for different temperatures $T$ (as shown in the legend), for stress along the armchair direction (left column) and the zigzag direction (right column), for increasing values of stress from top to bottom. The stress in the armchair direction is (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while along the zigzag direction is (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions given by Equation (23) are indicated by the solid curves.
4 Conclusions
We investigated the planar dynamics of a graphene sheet using Hamiltonian formalism and an efficient symplectic integration technique allowing the creation of accurate numerical data for very long simulation times. Our MD simulations examined the effects of finite temperatures in the mechanical response of graphene. In particular, we derived stress-strain responses for two different directions of applied stress, along either the armchair or the zigzag direction, at various temperatures.
We obtained a small, almost linear decrease of the Young’s modulus of graphene as the temperature of the sheet increases. Such a variation of Young’s modulus with temperature is in line with previous investigations. Furthermore, an intriguing temperature dependence has been obtained for the third-order elastic modulus, which is found to decrease (slightly increase) its absolute value with increasing temperature, for stresses along the armchair (zigzag) direction. Finally, we found that tensile strength and failure strain decrease approximately linearly with temperature and quantified the slope of this variation.
We also discussed the dependence of the distributions of bond lengths and bond angles within the graphene sheet, on both the applied stress and temperature. Approximate analytical expressions for these distributions are provided. In particular, we found that the distributions can be described by the sum of two Gaussian peaks, where the center of each peak is obtained from the values of bond lengths or bond angles, respectively, in the strained graphene subjected to the particular amount of stress at zero temperature. The variance of each peak as a function of temperature can be derived by the corresponding data at zero applied stress, while for the bond length distributions a scaling factor is additionally incorporated to account for the highly anharmonic nature of the Morse potential. Thus, we have provided a detailed description of the effects of both stress and temperature on the structural properties of graphene.
Acknowledgements: S. E. acknowledges support from the National Research Foundation of South Africa, the University of Cape Town (UCT), and the Science Faculty of UCT. All authors thank the Erasmus+/International Credit Mobility KA171 program for support. We thank the Centre for High Performance Computing (CHPC) of South Africa for providing the computational resources for this project. S. E. thanks M. Hillebrand, B. Many Manda, and A. Ngapasare for their insight regarding the numerical simulations, and D. Kruyt for their technical assistance with code optimization and systems administration.